OpenStudy (anonymous):

which is the simplified form of (x+4/2x+6)+(3/x^2-9)?

OpenStudy (unklerhaukus):

\[\frac{x+4}{2x+6}+\frac{3}{x^2-9}\]\[=\frac{x+4}{2(x+3)}+\frac{3}{(x+3)(x-3)}\]\[=\frac{(x+4)(x-3)}{2(x+3)(x-3)}+\frac{2\times3}{2(x+3)(x-3)}\]\[=\frac{(x+4)(x-3)+2\times3}{2(x+3)(x-3)}\]